OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by y=x^2, y=1, and the y-axis about the line y= -3

OpenStudy (anonymous):

is that a paraboloid ring?

OpenStudy (amistre64):



OpenStudy (amistre64):

its a washer method, simply becasue we are cutting out the middle; might as well call it the donut method

OpenStudy (amistre64):

lets add 3 to make life easier; y = x^2 +3; and y = 1+3 = 4

OpenStudy (anonymous):